OpenStudy (anonymous):

What is the value of the expression? 4^3 • 4^−5 A. 4^−15 B. 4^−2 C. 4^2 D. 4^8

OpenStudy (anonymous):

When bases are same and in Multiplication, their powers get added.. \[\large \color{green}{a^m \cdot a^n = a^{m+n}}\] Here, \(m = 3\) and \(n = -5\), And just plug in the values and do the necessary calculation thereafter.. :)

OpenStudy (lxelle):

So, 4^3+(-5)=???

OpenStudy (anonymous):

B

OpenStudy (lxelle):

Yeah

OpenStudy (haseeb96):